OpenStudy (anonymous):

Solve using the addition method. What is the y-value of the solution to the system? 2x – 3y = -1 3x + 4y = 24

OpenStudy (anonymous):

Do you mean elimination?

OpenStudy (anonymous):

addition method

OpenStudy (anonymous):

Anyways, Ill start it off. Addition method is another word for elimination method Choose to eliminate the variable that is easier. I choose to eliminate x. To do that I have to mult. the top by (3) and thw bottom by -2 2x-3y=-1 (2) 3x+4y=24 (-2) _______________________________ 6x-9y=-3 -6x-8y=-48 Add and eliminate the x ________________________________________ -17y= -51 y= \[\frac{ -51 }{ -17 }=\frac{ 51 }{ 17 }\]

OpenStudy (anonymous):

\[y=\frac{ 51}{ 17 }=3\]

OpenStudy (anonymous):

Since we have found y which is 3. Plug y (3) into an orginal equation. So ill post it at the top. 2x-3(3)=-1 2x-9=-1 +9 +9 ________________________________ 2x=8= ?????

OpenStudy (anonymous):

thank you soo much, u really helped me out on this one, thank you

OpenStudy (anonymous):

No problem. :)